# How to Purchase Labels? A Cost-effective Approach Using Active Learning Markets

### v2

@@ 3.3 Data purchase strategies
The data purchase problem addressed in this paper is inherently constrained by budget and governed by heterogeneous pricing. Active learning, a framework designed to select the most informative data point under label scarcity, provides a principled foundation for addressing such challenges (Zheng and Padmanabhan, 2006, 2002). In particular, active learning allows the analyst to determine which data points should be acquired to maximize model improvement while minimizing redundant or uninformative acquisitions. Active learning strategies are typically divided into two families: optimization-based and heuristic-based approaches (Cohn et al., 1996; Tong, 2001; Settles, 2011). Optimization-based methods rely on objective functions that formalize notions of informativeness, such as uncertainty reduction or expected loss. In contrast, heuristic-based approaches draw on approximations or model ensembles to identify uncertain points. In our study, we focus on one strategy from each category: variance-based active learning (VBAL), rooted in optimal experimental design, and Query-by-Committee active learning (QBCAL), a heuristic method based on committee disagreement. As benchmarks, we include RSC and a Greedy Knapsack (GK) baseline that greedily purchases the lowest-cost candidates by ranking them using $1/\eta_{j}$, since $l_{j}$ is not observable ex ante.
@@ 3.3.1 Variance-Based Active Learning (VBAL)
Variance-based active learning (VBAL) is based in Optimal Experimental Design (OED), which provides criteria for selecting data points that most effectively reduce model uncertainty in regression tasks (Lu et al., 2024; Atkinson, 1996; Atkinson and Bailey, 2001; Atkinson et al., 2007). A central quantity in OED is the information matrix $\mathbf{X}^{\top}\mathbf{X}$, derived from the feature $\mathbf{X}$ of the labelled dataset (Zheng and Padmanabhan, 2006). Several classical criteria operate on this matrix to guide selection: D-optimality seeks to minimize the volume of the confidence ellipsoid around the regression coefficients by maximizing $\det(\mathbf{X}^{\top}\mathbf{X})$; The A-optimality minimizes the average variance of the coefficient estimates by minimizing $\text{tr}(\mathbf{X}^{\top}\mathbf{X})^{-1}$; and the G-optimality focuses on predictive reliability, aiming to minimize the maximum prediction variance in the design space. While D and A-optimality optimize parameter estimation globally, G-optimality is more aligned with our goal: ensuring that predictions are uniformly reliable across all regions of the input space. In formal terms, G-optimality minimizes the worst-case prediction variance, defined as
@@ 3.3.2 Query-by-Committee-Based Active Learning (QBCAL)
Classical QBCAL approaches in classification typically fall into two categories: vote entropy (Freund et al., 1997; Argamon-Engelson and Dagan, 1999), which quantifies disagreement among classifiers by evaluating the entropy of predicted labels, and margin confidence (Abe and Mamitsuka, 1998; Settles, 2011), which prioritizes points with the smallest difference between the most probable class labels. Although effective in classification tasks, these strategies are not directly applicable in regression settings where the output is continuous rather than categorical.
@@ 4.1.1 Dataset and case-study
We used a public real estate data set from Sindian District, New Taipei City (Yeh, 2021), containing key transaction features (date, age, MRT distance, convenience stores) and unit-area price. A linear regression model is used to explain the variation in prices.
@@ 4.1.5 Is Our Active Learning Strategy Robust?
A Monte Carlo simulation with 1,000 replications was conducted by repeatedly resampling subsets of the original data, with a model-improvement threshold fixed at 20%. Figures 6 and 7 show the resulting distributions of effectively purchased data points (i.e., data points used to improve the model) for buyer-centric (BC) and seller-centric (SC) pricing schemes. VBAL and QBCAL exhibit concentrated frequency distributions (Figures 66(a)6(b) and 77(a)7(b)), indicating stable behaviour across replications, while the random (RSC) and the Greedy Knapsack (GK) baselines produce flatter distributions (Figures 66(c)6(d) and 77(c)7(d)), reflecting higher variability and lower robustness.
Parameter Sensitivity.
@@ 4.2.1 Case study and dataset
We use the BDG2 dataset (Wang et al., 2022), which contains the hourly energy consumption of 1,636 buildings from 2016 to 2017. For illustration, we consider an educational building (“Rachael”) as the data analyst and identify its most similar building (“Madge”) via Euclidean distance in the feature space, treating it as the data seller. Each observation includes the month, weekday, day, hour, holiday indicator, air temperature, and energy consumption. We adopt a sliding-window approach using only the “value” characteristic (hourly consumption) to construct lagged inputs: the consumption of the past seven days’ at 24-hour intervals (e.g. t168t{-}168, t144t{-}144, …, t24t{-}24), with the current hour’s consumption at time $t$ as the prediction target. Then a linear forecasting model is trained on these lagged features.

### v3

@@ 3.3 Data purchase strategies
The data purchase problem addressed in this paper is inherently constrained by budget and governed by heterogeneous pricing. Active learning, a framework designed to select the most informative data point under label scarcity, provides a principled foundation for addressing such challenges [43, 42]. In particular, active learning allows the analyst to determine which data points should be acquired to maximize model improvement while minimizing redundant or uninformative acquisitions. Active learning strategies are typically divided into two families: optimization-based and heuristic-based approaches [12, 36, 33]. Optimization-based methods rely on objective functions that formalize notions of informativeness, such as uncertainty reduction or expected loss. In contrast, heuristic-based approaches draw on approximations or model ensembles to identify uncertain points. In our study, we focus on one strategy from each category: variance-based active learning (VBAL), rooted in optimal experimental design, and Query-by-Committee active learning (QBCAL), a heuristic method based on committee disagreement. As benchmarks, we include RSC and a Greedy Knapsack (GK) baseline that greedily purchases the lowest-cost candidates by ranking them using $1/\eta_{j}$, since $l_{j}$ is not observable ex ante.
@@ 3.3.1 Variance-Based Active Learning (VBAL)
Variance-based active learning (VBAL) is based in Optimal Experimental Design (OED), which provides criteria for selecting data points that most effectively reduce model uncertainty in regression tasks [25, 7, 6, 8]. A central quantity in OED is the information matrix $\mathbf{X}^{\top}\mathbf{X}$, derived from the feature $\mathbf{X}$ of the labelled dataset [43]. Several classical criteria operate on this matrix to guide selection: D-optimality seeks to minimize the volume of the confidence ellipsoid around the regression coefficients by maximizing $\det(\mathbf{X}^{\top}\mathbf{X})$; The A-optimality minimizes the average variance of the coefficient estimates by minimizing $\text{tr}(\mathbf{X}^{\top}\mathbf{X})^{-1}$; and the G-optimality focuses on predictive reliability, aiming to minimize the maximum prediction variance in the design space. While D and A-optimality optimize parameter estimation globally, G-optimality is more aligned with our goal: ensuring that predictions are uniformly reliable across all regions of the input space. In formal terms, G-optimality minimizes the worst-case prediction variance, defined as
@@ 3.3.2 Query-by-Committee-Based Active Learning (QBCAL)
Classical QBCAL approaches in classification typically fall into two categories: vote entropy [18, 5], which quantifies disagreement among classifiers by evaluating the entropy of predicted labels, and margin confidence [2, 33], which prioritizes points with the smallest difference between the most probable class labels. Although effective in classification tasks, these strategies are not directly applicable in regression settings where the output is continuous rather than categorical.
@@ 4.1.1 Dataset and case-study
We used a public real estate data set from Sindian District, New Taipei City [40], containing key transaction features (date, age, MRT distance, convenience stores) and unit-area price. A linear regression model is used to explain the variation in prices.
@@ 4.1.5 Is Our Active Learning Strategy Robust?
A Monte Carlo simulation with 1,000 replications was conducted by repeatedly resampling subsets of the original data, with a model-improvement threshold fixed at 20%. Figures 6 and 7 show the resulting distributions of effectively purchased data points (i.e., data points used to improve the model) for buyer-centric (BC) and seller-centric (SC) pricing schemes. VBAL and QBCAL exhibit concentrated frequency distributions ((Figures 66(a)6(b) and 77(a)7(b)), indicating stable behaviour across replications, while the random (RSC) and the Greedy Knapsack (GK) baselines produce flatter distributions (Figures 66(c)6(d) and 77(c)7(d)), reflecting higher variability and lower robustness.
@@ 4.2.1 Case study and dataset
We use the BDG2 dataset [37], which contains the hourly energy consumption of 1,636 buildings from 2016 to 2017. For illustration, we consider an educational building (“Rachael”) as the data analyst and identify its most similar building (“Madge”) via Euclidean distance in the feature space, treating it as the data seller. Each observation includes the month, weekday, day, hour, holiday indicator, air temperature, and energy consumption. We adopt a sliding-window approach using only the “value” characteristic (hourly consumption) to construct lagged inputs: the consumption of the past seven days’ at 24-hour intervals (e.g. t168t{-}168, t144t{-}144, …, t24t{-}24), with the current hour’s consumption at time $t$ as the prediction target. Then a linear forecasting model is trained on these lagged features.